# Fair Contracts

### v1

@@ 1.2 Our Results
The hardness results in general settings motivate us to investigate the computational questions in more restrictive settings with a constant number of agents or a constant number of tasks. For the relatively easy case with a constant number of tasks, we first observe that the computation of the optimal envy-free contract can be solved in polynomial time via linear programming if the task assignment is exogenously given. With a constant number of tasks, the number of possible assignments is polynomial in the input size. Therefore, finding the optimal envy-free contract is polynomial-time solvable in this case (Proposition 4.1). This result also immediately implies efficient algorithms for finding optimal $\epsilon$-EF or EF1 contracts.
@@ 1.3 Related Work
Despite extensive research in algorithmic contract theory, contract design under fairness constraints has received limited attention. Fehr et al., (2007) conducted experimental studies showing that, when fairness concerns are present, bonus contracts can outperform standard incentive contracts. More recently, Dasaratha et al., (2023); Tłuczek et al., (2025) consider contract design with equity constraints. Their model is orthogonal to ours: they consider a single-task setting and restrict the compensation scheme to be identical across agents. In contrast, we study a multi-task environment, where the principal simultaneously assigns tasks and designs individualized incentive contracts to maximize profit, subject to envy-free constraints.
@@ 4.1 Polynomial-time Algorithms for a Constant Number of Tasks
The proof of Proposition 4.1 is in Appendix A.2. Note that this immediately implies poly-time algorithms for optimal $\epsilon$-EF and EF1 contracts (in Appendix B.3).
@@ 5 $\delta\leftarrow\frac{\epsilon}{m}$ ;
6 Define $\mathcal{D}_{i}$, $\mathcal{U}_{i}$, $\mathcal{H}$ as per Equation (4) and functions $\hat{U}_{i,j}(\cdot)$ and $\hat{U}^{P}(\cdot,\cdot)$ (Eq. (5);
@@ B.3 Computing EF1 and $\epsilon$-EF Contracts with Constant Number of Tasks
By proof of Proposition 4.1, the proof of $\epsilon$-EF is immediate. To prove it also holds for EF1 contracts, we note that given an allocation $S_{1},S_{2},\dots,S_{n}$, for each pair of agents $(i,j)$, the EF1 constraints can be represented as $|S_{j}|$ EF constraints by enumerating all task $g\in S_{j}$, i.e.,
$\sum_{k\in S_{i}}\alpha_{k}p_{i,k}r_{k}-c_{i,k}\geq\sum_{k\in S_{j}\backslash\{g\}}\max\{\alpha_{k}p_{i,k}r_{k}-c_{i,k},0\},\quad\forall j\neq i,\forall g\in S_{j}$
@@
Moreover, by introducing additional variables $t_{i,k}$ for all $i\in\mathcal{N},k\in\mathcal{M}$ and rewriting the program for EF1 contracts as in the proof of Proposition 4.1, we obtain a linear program with polynomial number of variables and constraints. This concludes the proof. ∎

### v2

@@ 1.2 Our Results
The hardness results in general settings motivate us to investigate the computational questions in more restrictive settings with a constant number of agents or a constant number of tasks. For the relatively easy case with a constant number of tasks, we first observe that the computation of the optimal envy-free contract can be solved in polynomial time via linear programming if the task assignment is exogenously given. With a constant number of tasks, the number of possible assignments is polynomial in the input size. Therefore, finding the optimal envy-free contract is polynomial-time solvable in this case (Proposition 4.1). Moreover, we design efficient algorithms for finding optimal $\epsilon$-EF or EF1 contracts using a similar framework but slightly involved analysis (Proposition B.3).
@@ 1.3 Related Work
Despite extensive research in algorithmic contract theory, contract design under fairness constraints has received limited attention. Fehr et al., (2007) conducted experimental studies showing that, when fairness concerns are present, bonus contracts can outperform standard incentive contracts. More recently, Dasaratha et al., (2023); Feng et al., (2024); Tłuczek et al., (2025) consider contract design with equity constraints. Their model is orthogonal to ours: they consider a single-task setting and restrict the compensation scheme to be identical across agents. In contrast, we study a multi-task environment, where the principal simultaneously assigns tasks and designs individualized incentive contracts to maximize profit, subject to envy-free constraints.
@@ 4.1 Polynomial-time Algorithms for a Constant Number of Tasks
The proof of Proposition 4.1 is in Appendix A.2. Relatedly, we show poly-time algorithms for optimal $\epsilon$-EF and EF1 contracts (in Appendix B.3), where the proof for EF1 contracts turns out to be slightly involved.
@@ 5 $\delta\leftarrow\frac{\epsilon}{m}$ ;
6 Define $\mathcal{D}_{i}$, $\mathcal{U}_{i}$, $\mathcal{H}$ as per Equation (4) and functions $\hat{U}_{i,j}(\cdot)$ and $\hat{U}^{P}(\cdot,\cdot)$ (Eq. (5));
@@ B.3 Computing EF1 and $\epsilon$-EF Contracts with Constant Number of Tasks
Following the proof of Proposition 4.1, the proof of $\epsilon$-EF is immediate.
The proof for EF1 contracts is more involved. We first note that given an allocation $S_{1},S_{2},\dots,S_{n}$, there are at most a constant number of agents assigned a non-empty set of tasks. The number of such allocations is at most $O(n^{m})$. Hence, as in the proof of Proposition 4.1, we can enumerate over all the allocations.
Hence, in the following we fix an allocation $S_{1},S_{2},\dots,S_{n}$. Consider the EF1 constraint relative to a pair of agents $(i,j)$, which guarantees that agent $i$ prefers $S_{i}$ over $S_{j}$ up to one item. Note that if we directly enumerate all the possible removable tasks $k_{i,j}$ (as in Definition 2.7) for all pairs of agents $(i,j)$, there would be an exponential number of possible sets of constraints, i.e., combinations of removal tasks. Our main result is that we actually only need to enumerate a polynomial number of possible sets of constraints. Then, for each allocation and combination of removal tasks, we solve a linear program.
For pair $(i,j)$, consider the following four cases:
  1. (i)
    $|S_{i}|>0$ and $|S_{j}|=0$
  2. (ii)
    $|S_{i}|=0$ and $|S_{j}|=0$
  3. (iii)
    $|S_{i}|>0$ and $|S_{j}|>0$
  4. (iv)
    $|S_{i}|=0$ and $|S_{j}|>0$
(i) By the IR constraints for agent $i$, the EF1 constraints obviously are satisfied.
(ii) Since both sides of EF1 constraints are 00, the EF1 constraints are satisfied.
(iii) There are at most $|S_{j}|$ different possible EF1 constraints, which are obtained by enumerating all tasks $k_{i,j}\in S_{j}$. Since there are at most $m$ agents such that $|S_{i}|>0$, we get that the total number of possible sets of EF1 constraints related to an agent $j$ is $O(m^{m})$. Moreover, there are at most $m$ such $S_{j}$, and hence the total number of possible sets is $O((m^{m})^{m})$.
(iv) Suppose the set of agents that are assigned the empty set is $A$. First, note that there are at most $n$ agents that are assigned empty sets of tasks. Hence, we cannot directly enumerate all the removable tasks $k_{i,j}$ for each agent $i$. However, the EF1 constraints of pairs $(i,j)$ imply that
$0\geq\sum_{k\in S_{j}\backslash\{k_{i,j}\}}\max\{\alpha_{k}p_{i,k}r_{k}-c_{i,k},0\},\quad\forall j\neq i,\exists k_{i,j}\in S_{j}$ (23)
which implies that for task $k\in S_{j}\setminus\{k_{i,j}\}$, the contract must be at most
$\alpha_{k}\leq w_{i,k}\triangleq\frac{c_{i,k}}{p_{i,k}r_{k}}$
This means that if the removable task is $k_{i,j}$, as long as $\alpha_{k}\leq w_{i,k}$ holds for all tasks $k\in S_{j}\setminus\{k_{i,j}\}$, the agent pair $(i,j)$ is EF1. Note that $w_{i,k}\triangleq\frac{c_{i,k}}{p_{i,k}r_{k}}$ is the minimum wage that is used to incentivize agent $i$ to exert effort for task $k$. Moreover, in this case, the contract $\alpha_{k_{i,j}}$ of the removable task $k_{i,j}$ has no restriction.
Given an EF1 contract, suppose the set of removable tasks from $S_{j}$ is $\{k_{i,j}\}_{i\in A}$ for the set of agents $A$. Note that $\{k_{i,j}\}_{i\in A}$ is a multiset. By the above discussion, it determines the upper bounds (i.e., the minimum wages) for all contracts $\alpha_{k}$ of all tasks $k\in S_{j}$, which is also a sufficient condition for EF1 constraints to hold for all agents $i\in A$ regarding set $S_{j}$. Hence, instead of enumerating removable tasks for all agents $i\in A$ regarding set $S_{j}$, we can equivalently enumerate the upper bounds for contracts of all tasks in $S_{j}$.
The algorithm for case (iv) works as in Algorithm 4.
5 Compute the minimum wages $w_{i,k}$ for all agents $i\in A$ and tasks $k\in S_{j}$ ;
6 Sort agents from low to high according to $\{w_{i,k}\}_{i\in A}$ for task $k$. Denote the list of agents $l_{k}$. ;
7 $\mathcal{F}\leftarrow\emptyset$ ;
8 for  every set of indexes $\{i_{k}\}_{k\in S_{j}}$ do
9       if $\{i_{k}\}_{k\in S_{j}}$ is feasible then
10             =$\mathcal{F}=\mathcal{F}\cup\{\{i_{k}\}_{k\in S_{j}}\}$ ;
11
12       end if
13
14 end for
return $\mathcal{F}$.
2 Input: Set of tasks $S_{j}$, set of unallocated agents $A$, $\{p_{i,j}\}_{i\in{A},j\in S_{j}}$, $\{c_{i,j}\}_{i\in{A},j\in{S_{j}}}$, $m,n$
Output: All feasible sets of indexes $\{i_{k}\}_{k\in S_{j}}$.
Algorithm 4 Find Upper Bounds for Contracts $\alpha_{k}$, $k\in S_{j}$
4 Input: Set of tasks $S_{j}$, set of unallocated agents $A$, $\{p_{i,j}\}_{i\in{A},j\in S_{j}}$, $\{c_{i,j}\}_{i\in{A},j\in{S_{j}}}$, $m,n$
Output: All feasible sets of indexes $\{i_{k}\}_{k\in S_{j}}$.
For each task $k\in S_{j}$, we calculate the minimum wage $w_{i,k}$ for agent $i\in A$. Then, for each task $k$, we sort all agents according to their minimum wages from low to high. Additionally, we add 00 to the head and $n+1$ to the tail, respectively, representing contract $\alpha_{k}=0$ and contract $\alpha_{k}=1$. Denote this list as $l_{k}$. Now, we have $|S_{j}|$ such lists. To determine the upper bounds for contract $\alpha_{k}$ of task $k\in S_{j}$, we can iterate over the ranked agents in list $l_{k}$ and find their corresponding minimum wages.
Suppose in one iteration, we obtain indexes $\{i_{k}\}_{k\in S_{j}}$ for all the lists. Then, given an allocation of sets $S_{1},S_{2},\dots,S_{n}$, we form a new program
$\max_{\{\alpha_{k}\}_{k\in\mathcal{M}}}\quad$ ${\rm Rev}(S,\alpha)$ (EF1-Program)
subject to (IR constraints)
EF1 constraints from case (iii)
Upper bounds $\alpha_{k}\leq w_{l_{k}[i_{k}],k}$ for all the tasks
However, not every enumeration of upper bounds is desired, as some of them may lead to a solution of (EF1-Program) that is not EF1. To see this, consider an agent $a\in A$ and two tasks $k,k^{\prime}\in S_{j}$. The indexes of agent $i$ in lists $l_{k},l_{k^{\prime}}$ are $a_{k},a_{k^{\prime}}$, respectively. Suppose in one enumeration, we have $i_{k}>a_{k}$ and $i_{k^{\prime}}>a_{k^{\prime}}$. By solving (EF1-Program), it may lead to one undesired solution where $w_{l_{k}[a_{k}],k}\leq\alpha_{k}\leq w_{l_{k}[i_{k}],k}$ and $w_{l_{k^{\prime}}[a_{k^{\prime}}],k^{\prime}}\leq\alpha_{k^{\prime}}\leq w_{l_{k^{\prime}}[i_{k^{\prime}}],k^{\prime}}$. Then, this solution is not EF1 for agent $a$.
To rule out the above undesired enumerated upper bounds, we need to check the feasibility of enumerated upper bounds as follows: If there exist no two lists $l_{k},l_{k^{\prime}}$ where $k,k^{\prime}\in S_{j}$ such that there is an agent $a\in A$ appearing in both $l_{k}[1,\dots,i_{k}-1]$ and $l_{k^{\prime}}[1,\dots,i_{k^{\prime}}-1]$, the enumeration of upper bounds is feasible.
Finally, we show that there are at most a polynomial number of possible sets of constraints in case (iv). Note that there are at most $O(n^{|S_{j}|})$ possible sets of upper bounds for the set $S_{j}$. Since there are at most $m$ tasks, the total number of possible sets of upper bounds is $O(n^{m})$.
@@ 7 $\mathcal{F}\leftarrow\emptyset$ ;
In summary of the above four cases, given a set allocation, we need to solve at most $O(m^{m^{2}}n^{m})$ different programs (EF1-Program). For each set of allocation, EF1 constraints, and upper bounds, we check if the enumerated upper bounds are feasible. In that case, we proceed to solve an LP. For each program (EF1-Program), by introducing additional variables $t_{i,k}$ for all $i\in\mathcal{N},k\in\mathcal{M}$ and rewriting EF1 contracts as in Proposition 4.1, we obtain a linear program with a polynomial number of variables and constraints. This concludes the proof. ∎